### v1

# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

Abstract
Risk assessment tools in healthcare commonly employ point-based scoring systems that map patients to ordinal risk categories via thresholds. While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) partial supervision arising from intervention-censored outcomes, where only extreme categories can be reliably labeled, and (2) asymmetric misclassification costs that increase with ordinal distance. We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds under these constraints. Our approach handles partial supervision through per-instance feasible label sets, incorporates asymmetric distance-aware objectives, and prevents middle-category collapse via minimum threshold gaps. We further develop a CSO relaxation using softplus losses that preserves the ordinal structure while enabling efficient optimization. The framework supports governance constraints including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows.
11footnotetext: Center for Health Systems and Policy Modeling, Department of Health Policy and Management, Johns Hopkins University, Baltimore, MD, USA. Email: fganjkh1@jhu.edu22footnotetext: Malone Center for Engineering in Healthcare, Johns Hopkins University, Baltimore, MD, USA33footnotetext: Center for Systems Science and Engineering, Department of Civil and Systems Engineering, Johns Hopkins University, Baltimore, MD, USA. Emails: espring6@jh.edu, kimia@jhu.edu44footnotetext: Department of Physical Medicine and Rehabilitation, School of Medicine, Johns Hopkins University, Baltimore, MD, USA. Email: ehoyer1@jhmi.edu55footnotetext: Johns Hopkins Hospital, Baltimore, MD, USA66footnotetext: Department of Physical Therapy, University of Nevada, Las Vegas, Las Vegas, NV, USA. Email: daniel.young@unlv.edu
1 Introduction
Across healthcare settings, risk assessment is routinely implemented via itemized, checklist-style instruments in which clinicians mark the presence, absence, or severity of predefined risk factors. Each factor contributes a predetermined point value to a total score, and pre-specified thresholds on this score map patients to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [19, 16, 11, 3, 1]. This linear scoring combined with threshold-based categorization paradigm is widely adopted due to its transparency, auditability, and ease of operationalization [23, 25, 10]. The Johns Hopkins Fall Risk Assessment Tool (JHFRAT) exemplifies this approach, evaluating eight risk factors including age, fall history, medications, mobility, and cognition [19]. Each factor is assigned integer points (e.g., age $\geq$ 80 years: 3 points; impaired mobility: 2 points), and the total score maps to risk categories via thresholds: Low (0-5 points), Medium (6-13 points), and High ($\geq$ 14 points). These categories trigger standardized intervention bundles ranging from universal precautions to intensive monitoring and physical therapy consultations [17, 6].
The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating point weights, adjusting thresholds, and encoding deployment constraints while preserving the interpretable structure that clinicians trust [9, 20]. However, the practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches. The first challenge is partial supervision with selective labels. In deployment, patient outcomes and clinical interventions are inherently coupled. For fall risk assessment, we can construct high-confidence labels for extremes: patients who experience falls despite interventions (Safe-High) and patients who remain fall-free under minimal intervention (Safe-Low). However, patients receiving intermediate interventions have censored counterfactual outcomes–we cannot observe what would have occurred without intervention. This creates systematic label uncertainty for middle categories, necessitating a partial-label learning framework where each instance $i$ has a feasible set $\mathcal{S}_{i}$ rather than a single label. The second challenge is asymmetric, distance-aware misclassification costs. Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events, while over-triage (classifying a low-risk patient as high-risk) merely increases resource utilization. Moreover, the severity of misclassification increases with ordinal distance: misclassifying a high-risk patient as low-risk is more consequential than misclassifying them as medium-risk. Training objectives must therefore incorporate both directional asymmetry and distance-awareness aligned with clinical priorities [7].
Definition 1 (Ordinal Risk Assessment Problem).
Given training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $\mathcal{S}_{i}\subseteq\mathcal{K}=\{1,\ldots,K\}$ are feasible label sets indicating possible risk categories, and $w_{i}>0$ are instance weights, the goal is to learn parameters $(\beta,\tau)$ where $\beta\in\mathbb{R}^{p}$ defines a linear scoring function $s(x)=\beta^{\top}x$ and $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ are monotone thresholds, to minimize expected asymmetric ordinal loss subject to interpretability and operational constraints.
While ordinal regression methods have been extensively studied, existing approaches fail to address the unique challenges of clinical risk assessment. Classical ordinal regression methods such as the proportional odds model [14] assume parallel decision boundaries and complete supervision across all categories. Support vector approaches for ordinal regression [4] and threshold models [24] similarly require fully labeled training data. Recent extensions to deep learning for ordinal problems [15, 2] and specialized ranking losses [22] maintain these restrictive assumptions, treating all categories as equally important and all misclassification costs as symmetric. The partial label learning literature addresses scenarios where each instance is associated with a set of candidate labels [5, 27], with disambiguation methods that iteratively refine label assignments [28, 26] and loss-based methods that modify training objectives [8, 13]. However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) we require preservation of ordinal structure throughout the learning process, and (3) misclassification costs are inherently asymmetric and distance-aware in healthcare applications [12]. Furthermore, while extensions like partial proportional odds models [18] relax the proportional odds assumption, they still assume complete supervision and do not address the selective labeling problem where entire categories may lack reliable labels.
We introduce a constrained mixed-integer optimization framework that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption. Our approach formalizes the partial-label learning problem for ordinal risk assessment, where each instance $i$ has a feasible label set $\mathcal{S}_{i}$ encoding label uncertainty, and the optimization respects these constraints without fabricating labels for ambiguous cases. We jointly optimize $(\beta,\tau)$ to capture their interdependence under the ordinal structure. We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities. When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification. Finally, we derive a constrained score optimization (CSO) relaxation using softplus losses that preserves the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems, extending ideas from smooth hinge losses [21] to the ordinal setting with partial supervision.
The remainder of this paper is organized as follows. Section 2 presents our mixed-integer programming formulation and convex relaxation. Section 3 describes experimental evaluation on clinical risk assessment tasks. Section 4 discusses computational considerations, limitations, and deployment aspects. Section 5 concludes with implications for precision medicine.
2 Methods
2.1 Problem Formulation and Notation
We consider the problem of learning a risk stratification tool that maps patient features $x\in\mathbb{R}^{p}$ to ordinal risk categories $\mathcal{K}=\{1,2,\ldots,K\}$ (e.g., Low, Medium, High for $K=3$). The tool consists of a linear scoring function $s(x)=\beta^{\top}x$ and monotone thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ that partition the score space into $K$ categories. A patient with score $s(x)$ is assigned to category $k$ if $\tau_{k-1}<s(x)\leq\tau_{k}$, where we define $\tau_{0}=$-$\infty$ and $\tau_{K}=+\infty$ for notational convenience.
The key innovation in our formulation is handling partial supervision through feasible label sets. For each training instance $i$, rather than observing a single true label $y_{i}\in\mathcal{K}$, we have a feasible set $\mathcal{S}_{i}\subseteq\mathcal{K}$ that contains the possible true categories. This formulation naturally captures the selective labeling problem in clinical settings: we can confidently identify extreme cases ($\mathcal{S}_{i}=\{1\}$ for definitely low-risk, $\mathcal{S}_{i}=\{K\}$ for definitely high-risk) while middle-risk patients who received preventive interventions have ambiguous labels ($\mathcal{S}_{i}=\{1,2\}$ or $\mathcal{S}_{i}=\{2,3\}$ or even $\mathcal{S}_{i}=\{1,2,3\}$).
2.2 Mixed-Integer Programming Formulation
We formulate the joint learning of scoring weights $\beta\in\mathbb{R}^{p}$ and thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})$ as a mixed-integer program that handles partial supervision and asymmetric costs. The MIP approach provides exact solutions and naturally incorporates discrete constraints essential for clinical deployment.
Table 1 summarizes the decision variables and their interpretations. The key insight is to use binary indicators $y_{ik}$ to track whether instance $i$’s score exceeds each threshold $\tau_{k}$, and assignment variables $z_{ik}$ to determine the final category.
Table 1: Decision variables and parameters in the MIP formulation
Symbol Domain Description
$\beta$ $\mathbb{R}^{p}$ Scoring weights for features
$\tau_{k}$ $\mathbb{R}$ Threshold between categories $k$ and $k+1$
$s_{i}$ $\mathbb{R}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
$y_{ik}$ $\{0,1\}$ Indicator: score $s_{i}$ exceeds threshold $\tau_{k}$
$z_{ik}$ $\{0,1\}$ Assignment of instance $i$ to category $k$
$M$ $\mathbb{R}_{+}$ Big-M constant for constraint activation
$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter for numerical stability
$\delta$ $\mathbb{R}_{+}$ Minimum gap between consecutive thresholds
The complete MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{K}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (1a)
subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i$ (1b)
$\tau_{k}\leq\tau_{k+1}-\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (1c)
$s_{i}-\tau_{k}\geq\varepsilon-M(1-y_{ik}),\quad\forall i,k$ (1d)
$s_{i}-\tau_{k}\leq-\varepsilon+My_{ik},\quad\forall i,k$ (1e)
$z_{i1}\leq 1-y_{i1},\quad\forall i$ (1f)
$z_{ik}\leq y_{i,k-1}-y_{ik},\quad\forall i,k\in\{2,\ldots,K-1\}$ (1g)
$z_{iK}\leq y_{i,K-1},\quad\forall i$ (1h)
$\beta\in\Omega,\quad y_{ik},z_{ik}\in\{0,1\}$ (1i)
Equation (1b) ensures each instance is assigned to exactly one category. Constraint (1c) maintains threshold ordering with minimum separation $\delta$ to prevent degeneracy. Constraints (1d)-(1e) implement the big-M formulation linking scores to threshold crossings: $y_{ik}=1$ if and only if $s_{i}\geq\tau_{k}+\varepsilon$. The margin $\varepsilon>0$ ensures numerical stability and prevents instances from lying exactly on boundaries. Constraints (1f)-(1h) encode the logical relationship between threshold crossings and category assignments: instance $i$ is in category $k$ if it exceeds exactly the first k1k-1 thresholds.
The cost function $c(k,\mathcal{S}_{i})$ in objective (1a) encodes both the distance to the feasible set and directional asymmetry:
$c(k,\mathcal{S}_{i})=\min_{k^{*}\in\mathcal{S}_{i}}\ell(k,k^{*})$ (2)
This formulation assigns to each instance the cost of predicting category $k$ when the true category is the closest feasible one. For singleton feasible sets ($|\mathcal{S}_{i}|=1$), this reduces to standard supervised learning. For larger feasible sets, the optimization naturally selects the most consistent label within the feasible set.
We use an asymmetric ordinal loss that captures clinical priorities:
$\ell(k,k^{*})=\begin{cases}\alpha_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\text{ (under-triage)}\\ \alpha_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\text{ (over-triage)}\\ 0&\text{if }k=k^{*}\end{cases},$ (3)
with $\alpha_{\text{under}}>\alpha_{\text{over}}$ to penalize under-triage more heavily. The exponent $q\geq 1$ controls the growth rate: $q=1$ yields linear penalty growth with ordinal distance, while $q=2$ imposes quadratic penalties for severe misclassifications. In fall risk assessment, we typically set $\alpha_{\text{under}}/\alpha_{\text{over}}\approx 3$ to reflect that missing a high-risk patient is approximately three times worse than over-treating a low-risk patient.
The big-M constraints in (1d)-(1e) require careful selection of $M$ to ensure correctness while maintaining numerical stability. We set $M=2\cdot\max_{i}\|x_{i}\|\cdot B_{\beta}+B_{\tau}$ where $B_{\beta}$ is an upper bound on $\|\beta\|$ and $B_{\tau}$ bounds the threshold range. These bounds can be derived from domain knowledge (e.g., reasonable score ranges in clinical tools) or data-driven estimates. The MIP has $O(nK)$ binary variables and $O(nK)$ constraints, making it tractable for moderate-sized clinical datasets (thousands of patients) using modern solvers. For larger problems, we employ the CSO warm-start strategy described in Section 2.6.
2.3 Non-Degenerate Categories
When middle categories lack labeled examples (common when $\mathcal{S}_{i}\in\{\{1\},\{K\}\}$ for most instances), the optimization may collapse thresholds to eliminate these categories. While this might optimize the training objective, it defeats the clinical purpose of having gradated risk levels for differential intervention. To prevent this while avoiding label fabrication, we impose minimum gap constraints.
For the three-category case (Low/Medium/High), we enforce:
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min},$ (4)
where $\Delta_{\min}>0$ ensures the medium category maintains meaningful width. This parameter can be set through several approaches:
  1. 1.
    Incumbent-based: Use the gap from existing validated tools (e.g., JHFRAT has $\Delta_{\min}=8$ points between Low and High thresholds)
  2. 2.
    Cross-validation: On a held-out set with known middle-category examples, select $\Delta_{\min}$ that maximizes classification performance
  3. 3.
    Clinical reasoning: Set based on meaningful score differences (e.g., requiring at least 2-3 risk factors to differ between Low and High)
For $K>3$ categories, we can impose similar constraints on consecutive gaps or the total range, depending on which categories lack supervision.
2.4 Governance and Interpretability Constraints
The feasible set $\Omega$ in Model (1i) encodes interpretability and deployment requirements essential for clinical adoption, in particular:
Sign constraints encode clinical knowledge about risk factor directionality. For factors known to increase risk (e.g., age $\geq$ 80, fall history), we enforce $\beta_{j}\geq 0$. For protective factors (e.g., independent mobility), we enforce $\beta_{j}\leq 0$. This prevents counterintuitive weights that would undermine clinical trust:
$\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},\quad\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$ (5)
Sparsity constraints limit cognitive load by bounding the number of active features. Clinical tools typically use 5-10 items for practical usability:
$\sum_{j=1}^{p}u_{j}\leq s_{\max},\quad-Mu_{j}\leq\beta_{j}\leq Mu_{j},\quad u_{j}\in\{0,1\},$ (6)
where binary variables $u_{j}$ indicate whether feature $j$ is active. The constraint ensures that at most $s_{\max}$ features have non-zero weights.
Minimal modification constraints facilitate adoption by limiting changes from incumbent tools. Clinicians are more likely to accept tools that refine rather than replace existing practice:
$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|,$ (7)
where $\beta^{(0)}$ represents the incumbent weights. The first form enforces hard limits on individual changes, while the second adds a soft penalty to the objective.
Grouping constraints ensure related features are treated consistently. For example, different mobility assessments might be grouped:
$\beta_{j}=\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G},$ (8)
where $\mathcal{G}$ contains pairs of features that should have identical weights.
Performance constraints ensure the optimized model meets minimum clinical safety requirements. Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly bound error rates. For instances with known true labels (singleton feasible sets $\mathcal{S}_{i}=\{k_{i}^{*}\}$), we define:
False positive rate constraint (e.g., limiting low-risk patients classified as high-risk):
$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\alpha_{\text{FP}}.$ (9)
False negative rate constraint (e.g., limiting high-risk patients classified as low-risk):
$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\alpha_{\text{FN}}.$ (10)
More generally, for any true category $k^{*}$ and predicted category $k$, we can constrain:
$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq\alpha_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq\gamma_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (11)
where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ is the set of instances with true label $k^{*}$, and $\alpha_{k^{*},k}$ (or $\gamma_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
These constraints are linear in the assignment variables $z_{ik}$ and integrate directly into the MIP formulation (19). They are particularly valuable for ensuring that optimization does not sacrifice critical safety metrics (e.g., sensitivity for high-risk patients) in favor of overall accuracy. Note that these constraints only apply to instances with known ground truth (singleton feasible sets), which aligns with the partial supervision framework.
2.5 Constrained Score Optimization (CSO) Relaxation
The MIP formulation provides exact solutions but faces computational challenges for large datasets. We derive a convex relaxation that preserves the essential structure—ordinal relationships, partial supervision, and asymmetric costs—while enabling efficient optimization via gradient-based methods.
The MIP model uses binary variables to encode category assignments and threshold crossings. The CSO relaxation replaces these discrete decisions with continuous margin-based losses. An important insight is that the ordinal loss decomposes into a sum of boundary crossing penalties, which we can approximate smoothly. Consider predicting category $k$ for an instance with true category $k^{*}$. The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$. We can rewrite this as:
  • If $k<k^{*}$: penalty for failing to cross thresholds $\tau_{k},\tau_{k+1},\ldots,\tau_{k^{*}-1}$.
  • If $k>k^{*}$: penalty for incorrectly crossing thresholds $\tau_{k^{*}},\tau_{k^{*}+1},\ldots,\tau_{k-1}$.
This decomposition motivates a margin-based formulation where we penalize violations of desired threshold crossings.
To create a smooth approximation, we replace discrete boundary indicators with smooth softplus penalties. For each instance-threshold pair $(i,k)$, we define
$\phi_{i,k}^{-}$ $=\frac{1}{\alpha}\log\left(1+\exp(\alpha(\tau_{k}-s_{i}))\right)$ (12)
[…]
Similarly, $\phi_{i,k}^{+}=\frac{1}{\alpha}\log(1+\exp(\alpha(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we shouldn’t. The temperature parameter $\alpha$ controls the approximation quality: as $\alpha\to\infty$, softplus approaches the hinge loss; as $\alpha\to 0$, it becomes linear.
To derive the CSO objective for instance $i$ with feasible set $\mathcal{S}_{i}$, we define the CSO loss function as:
[…]
The weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. To reproduce the asymmetric ordinal loss (3), we set:
$\lambda_{j}^{-}$ $=\alpha_{\text{under}}\cdot w_{j}\quad\text{(weight for failing to exceed threshold }j\text{)},$ (14)
$\lambda_{j}^{+}$ $=\alpha_{\text{over}}\cdot w_{j}\quad\text{(weight for incorrectly exceeding threshold }j\text{)}.$ (15)
where $w_{j}$ encodes the positional importance (e.g., $w_{j}=1$ for linear growth, $w_{j}=j$ for quadratic).
The minimization over $k\in\mathcal{S}_{i}$ implements partial supervision: the loss encourages the instance toward the most compatible category within its feasible set. This is convex since it’s the minimum of convex functions over a finite set.
The full CSO formulation can be written as follows.
$\min_{\beta,\tau}\quad$ $\sum_{i=1}^{n}w_{i}\mathcal{L}_{i}(\beta,\tau)+\mu R(\beta)$ (16a)
subject to $\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{K-1}$ (16b)
$\tau_{k+1}-\tau_{k}\geq\delta,\quad\forall k\in\{1,\ldots,K-2\}$ (16c)
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min}\quad\text{(for $K=3$)}$ (16d)
$\beta\in\Omega$ (16e)
where $R(\beta)$ is an optional regularizer (e.g., $\|\beta\|_{2}^{2}$ for ridge, $\|\beta\|_{1}$ for lasso via proximal methods), and $\Omega$ encodes the same governance constraints as the MIP.
The CSO objective is differentiable with respect to both $\beta$ and $\tau$, with gradients
$\frac{\partial\mathcal{L}_{i}}{\partial\beta}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\left[\sum_{j=1}^{k-1}\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))-\sum_{j=k}^{K-1}\lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))\right]x_{i},$ (17)
$\frac{\partial\mathcal{L}_{i}}{\partial\tau_{j}}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\begin{cases}-\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))&\text{if }j<k\\ \lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))&\text{if }j\geq k\end{cases},$ (18)
where $\sigma(t)=1/(1+\exp(-t))$ is the sigmoid function, and $\mathcal{S}_{i}^{*}\subseteq\mathcal{S}_{i}$ contains the categories achieving the minimum in (13). When $|\mathcal{S}_{i}^{*}|>1$, we use subgradients or smooth approximations. The projection operators maintain feasibility:
  • $\Pi_{\Omega}(\beta)$: projects onto governance constraints (e.g., sign, sparsity)
  • $\Pi_{\mathcal{T}}(\tau)$: projects onto ordered thresholds with minimum gaps via isotonic regression with spacing constraints
Proposition 1 (Convexity).
The CSO objective (16) is convex in $(\beta,\tau)$ jointly.
Proof sketch.
The softplus functions $\phi_{i,k}^{\pm}$ are convex. The weighted sum in (13) preserves convexity. The minimum over $k\in\mathcal{S}_{i}$ is convex as the pointwise minimum of convex functions. The constraints form a convex set. ∎
Proposition 2 (Approximation Quality).
As $\alpha\to\infty$, the CSO solution approaches the MIP solution on separable data.
Proposition 2 follows from the softplus converging to the hinge loss, which exactly captures threshold crossing violations.
Algorithm 1 CSO-Guided Mixed-Integer Optimization for Risk Scoring
1:Training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$, parameters $\delta,\Delta_{\min},\alpha$, convergence tolerance $\epsilon$
2:Scoring weights $\beta^{*}$, thresholds $\tau^{*}$
3:Phase 1: CSO Warm-Start
4:Initialize $\beta^{(0)},\tau^{(0)}$ from incumbent tool or random
5:for $t=1,2,\ldots$ until convergence do
6:  Compute CSO loss gradients via (13)
7:  $\beta^{(t+1)}\leftarrow\Pi_{\Omega}(\beta^{(t)}-\eta_{t}\nabla_{\beta}\mathcal{L})$
8:  $\tau^{(t+1)}\leftarrow\Pi_{\mathcal{T}}(\tau^{(t)}-\eta_{t}\nabla_{\tau}\mathcal{L})$ $\triangleright$ Maintain ordering and gaps
9:  if $\|\beta^{(t+1)}-\beta^{(t)}\|+\|\tau^{(t+1)}-\tau^{(t)}\|<\epsilon$ then
10:   break
11:  end if
12:end for
[…]
14:Phase 2: MIP Refinement
15:Initialize MIP solver with warm-start solution $(\beta_{\text{warm}},\tau_{\text{warm}})$
16:Set initial bounds: $\beta_{j}\in[\beta_{\text{warm},j}-\rho,\beta_{\text{warm},j}+\rho]$
17:Solve MIP (19) with warm-start and bounds
18:Optional: If integer scores required, add $\beta_{j}\in\mathbb{Z}$ constraints
19:return optimal solution $(\beta^{*},\tau^{*})$
2.6 Two-Phase Optimization: CSO Warm-Start with MIP Refinement
Algorithm 1 presents our two-phase optimization approach. The CSO relaxation in Phase 1 provides a high-quality warm-start solution that significantly accelerates the subsequent MIP solve in Phase 2. This hybrid approach combines the computational efficiency of convex optimization with the exactness of mixed-integer programming. The warm-start not only reduces solve time but also helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
While our formulation allows continuous-valued scoring weights $\beta\in\mathbb{R}^{p}$, practical deployment often benefits from integer-valued scores that are easier for clinicians to compute mentally. Following the approach in [25], we can enforce integrality by adding constraints $\beta_{j}\in\mathbb{Z}$ to the MIP formulation. When combined with appropriate scaling of features, this yields point values similar to existing tools (e.g., 0-5 points per item). The trade-off between interpretability and predictive performance can be explicitly controlled through the granularity of allowed integer values. Our CSO warm-start remains valuable even with integer constraints, as rounding the continuous solution provides an excellent initial integer feasible solution.
3 Experiments
The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: partial supervision due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually. In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\mathbb{R}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$. The scoring function $s(x_{i})=\beta^{\top}x_{i}$ computes the total JHFRAT score, where $\beta\in\mathbb{Z}^{17}_{+}$ contains the integer point values for each risk factor.
The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk. Following our framework, we construct feasible label sets $\mathcal{S}_{i}$ based on observed outcomes and intervention patterns. Patients who fell despite interventions are confidently labeled as high-risk ($\mathcal{S}_{i}=\{3\}$), while those who remained fall-free without targeted interventions are labeled as low-risk ($\mathcal{S}_{i}=\{1\}$). The remaining patients who received interventions but did not fall have uncertain risk—they might be truly high-risk patients successfully protected by interventions, or lower-risk patients who received unnecessary precautions. Rather than forcing labels on these ambiguous cases, our framework excludes them from the primary optimization while using them for post-hoc validation.
We evaluate our framework on optimizing the Johns Hopkins Fall Risk Assessment Tool using electronic health record data from three Johns Hopkins Health System hospitals: Johns Hopkins Hospital, Johns Hopkins Bayview Medical Center, and Howard County Medical Center. Our dataset comprises 54,209 hospital admissions between March 28, 2022 and October 27, 2023, with complete JHFRAT assessments, intervention records, and fall events. The JHFRAT assessment consists of 17 non-zero coefficient risk factors across seven categories: age, bladder/bowel elimination, cognition, fall history, patient care equipment, medications, and mobility. On average, patients are assessed twice daily with JHFRAT, and our dataset has an average of 1.87 JHFRAT records per patient day.
3.1 Dataset Construction and Partial Labels
A total of 498 hospital encounters in the dataset include at least one fall event, constituting 0.92% of encounters and equating to 1.07 falls per 1,000 patient-days. The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach. The majority of patients in the data receive one or more fall-prevention interventions during hospitalization. Of the 31 possible fall-prevention interventions, 13 were identified by a team of clinicians as “targeted” for being resource-intensive and capable of meaningfully altering fall risk. Such targeted interventions include the use of bed exit alarms, increased patient rounding, and constant monitoring. We assume that these targeted interventions, due to their intensive nature and established efficacy, can meaningfully obscure underlying fall risk. In contrast, we do not consider general interventions, like education to patients and family about fall risks, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence.
Figure 1 illustrates how we partition the 54,209 patient encounters into three cohorts based on fall events and targeted intervention receipt. The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls has different interpretations depending on intervention status. We partition patient encounters into three cohorts with distinct labeling strategies:
  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled Safe High-Risk.
  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled Safe Low-Risk.
  • Intervention cohort: Encounters with at least one targeted intervention during the encounter, and no fall events. These encounters are considered to have uncertain risk, and are not labeled for the optimization.
Figure 1: Stratification of inpatient admissions based on fall events and targeted interventions. The intervention cohort was excluded from primary optimization analyses but retained for post-optimization validation.
The extreme imbalance in cohort sizes–with 80.7% of patients in the ambiguous intervention cohort–underscores why standard supervised learning approaches fail in this setting. Simply assigning these 43,799 patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data. Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 10,410 patients (19.3%) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.
3.2 MIP Models
The JHFRAT scale is divided, per current practice, into three risk categories: Low-Risk ($k=1$) for scores less than 6, Moderate-Risk ($k=2$) for scores 6-13, and High-Risk ($k=3$) for scores greater than 13. While our general framework jointly optimizes both scoring weights $\beta$ and thresholds $\tau$, the JHFRAT application presents unique deployment constraints that motivate a modified approach. To maintain compatibility with existing clinical workflows, electronic health record systems, and staff training, we fix the thresholds at their current values ($\tau_{1}=6,\tau_{2}=13$) and focus optimization on the scoring weights $\beta$ alone. This decision reflects a common practical scenario where healthcare systems require incremental refinements to existing tools rather than complete redesigns, as changing risk thresholds would necessitate updates to intervention protocols, clinical guidelines, and quality metrics across the entire health system.
The absence of reliable moderate-risk labels in our dataset further supports this constrained optimization approach. Due to the selective labeling problem, we can confidently identify only extreme cases: patients who fell despite interventions (Safe High-Risk) and those who remained fall-free without targeted interventions (Safe Low-Risk). Patients with moderate JHFRAT scores predominantly received interventions (78.3% in our data), making their true risk indeterminate—they might be correctly classified moderate-risk patients, or they might be misclassified high or low-risk patients. Rather than introducing noise by guessing labels for this ambiguous middle group, we apply singleton feasible sets only to the extreme cases: $\mathcal{S}_{i}=\{1\}$ for all $i$ in the Safe Low-Risk cohort and $\mathcal{S}_{i}=\{3\}$ for all $i$ in the Safe High-Risk cohort.
This configuration represents a special but important case of our general framework where partial supervision is extreme—no labeled examples exist for intermediate categories. The fixed thresholds naturally prevent the middle category collapse problem described in Section 2.3, as the moderate-risk band maintains its width of $\tau_{2}-\tau_{1}=7$ points regardless of the optimization. While this approach doesn’t leverage our framework’s full capability for joint score-threshold learning, it demonstrates its flexibility in handling severely limited supervision scenarios common in healthcare, where intervention effects obscure risk for the majority of patients. Future work could explore threshold optimization using external validation data or prospective trials where moderate-risk labels can be more reliably determined.
We apply three types of governance constraints in our MIP formulations for JHFRAT. Firstly, we apply a uniform lower and upper bound on all model coefficients. Since we only include the risk factors from JHFRAT with positive coefficients, it is reasonable to impose the clinical assumption inherent in the lower bound of 0 that all of these factors can only contribute positively to high fall risk. Furthermore, we limit coefficients to the highest class’s lower bound of 13 for model stability. Secondly, we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment. Finally, we impose monotonicity constraints for hierarchical risk factors to mantain clinical validity. In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e. the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs).
The complete JHFRAT-optimizing MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (19a)
subject to $\sum_{k=1}^{3}z_{ik}=1,\quad\forall i$ (19b)
(1d) - (1h) (19c)
$0\leq\beta_{j}\leq 13\quad\forall j$ (19d)
$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (19e)
$\beta_{j}\in\mathbb{Z}\quad y_{ik},z_{ik}\in\{0,1\}$ (19f)
We consider four variations to the MIP model above:
  1. 1.
    Symmetric: The ordinal loss for under-triage and over-triage is the same ($\alpha_{under}=\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  2. 2.
    Asymmetric: The ordinal loss for under-triage is three times that of over-triage ($\alpha_{under}=3,\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  3. 3.
    False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). Let $\bar{\beta}$ indicate the JHFRAT model coefficients and let $\bar{z}$ indicate the feasible set binary classification variables imposed deterministically by the MIP formulation with $\bar{\beta}$. We impose the following two constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i1}$ (20a)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i2}$ (20b)
    These additional performance constraints ensure that the number of Safe Low-Risk labeled encounters classified as Moderate-Risk (20a) and High-Risk (20b) cannot exceed the numbers of such classifications with the JHFRAT model.
  4. 4.
    False-Negative (FN) & False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage. The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). With the same definition of $\bar{z}$ as above, we impose the following four constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|2\in S_{i}}z_{i0}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i0}$ (21a)
    $\sum_{i\in[n]|2\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i1}$ (21b)
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq 0.05\sum_{i\in[n]|0\in S_{i}}1$ (21c)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq 0.3\sum_{i\in[n]|0\in S_{i}}1$ (21d)
    The number of Safe High-Risk labeled encounters classified as Low-Risk (21a) and Moderate Risk (21b) cannot exceed the numbers of such classifications with the JHFRAT model. The portion of Safe Low-Risk labeled encounters classified as Moderate-Risk (21c) and High-Risk (21d) cannot exceed given thresholds (30% and 5% respectively). These constraints allow for a small increase from the JHFRAT benchmark of 26.5% and 0.9% for these same categorizations, as seen in Table 2.
We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability. The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign.
3.3 Experimental Setup
We split the encounters from the combined fall and reference cohorts into an 80% (encounters) training and validation set, and a 20% (encounters) test set, stratified by risk label. We retain the intervention cohort separately as an exogenous dataset for concordance analysis post-optimization. We perform 5-fold cross-validation for training the CSO, and run a hyperparameter gridsearch over the values of $\alpha_{under}$, $\alpha_{over}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$). For each set of hyperparameters, we consider the average coefficients across the cross-validation folds. Each MIP model is trained on the entire training set, utilizing the best feasible rounded coefficients from the CSO gridsearch. If none of the rounded CSO coefficients are feasible for a given MIP model, the JHFRAT coefficients are used for the warm-start.The CSO model is solved with cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
3.4 Baselines and Evaluation Metrics
We compare the outcomes of our approach against the original JHFRAT model with expert-derived weights. Furthermore, we compare the CSO performance with those of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline. We utilize the following metrics to compare model performance:
  • Classification Accuracy: We define tight accuracy as the portion of encounters correctly classified into low or high-risk categories. We further define loose accuracy to additionally include all encounters classified as moderate-risk as correctly classified.
  • High-Risk Precision and Recall: To account for the imbalance of the dataset, we rely on the high-risk precision and recall in addition to the accuracy metrics.
  • Binary Classification Performance Measures: We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves. We then utilize the area under each curve, AUROC and AUPRC respectively, as model performance metrics.
4 Results
4.1 Optimized Risk Factor Coefficients
The optimal solutions to the four MIP variations are featured in Table 2, in comparison with the current JHFRAT scoring coefficients. The optimal CSO coefficients for symmetric class loss (hyperparameters $\alpha_{under},\alpha_{over}=1,\lambda_{1}^{-},\lambda_{1}^{+},\lambda_{2}^{-},\lambda_{2}^{+}=1$) are also included as a representative example of CSO results.
Table 2: Risk factor coefficients across baseline and optimized models
Risk Factor JHFRAT CSO Symmetric Asymmetric FP-cap FN&FP-cap
Age 60-69 years 1 1.70 1 0 1 0
Age 70-79 years 2 1.92 2 3 1 1
Age $\geq$ 80 years 3 2.22 2 13 4 2
Incontinence 2 7.33 11 9 4 12
Urgency/frequency 2 3.75 3 6 5 3
Altered awareness 1 7.66 12 9 12 10
Impulsive 2 15.01 13 11 12 7
Lack of understanding 4 5.20 13 12 3 12
One present 1 0.77 0 3 2 2
Two present 2 1.38 1 3 3 6
Three or more present 3 1.56 2 13 3 8
Fall within 6 months 5 2.22 0 6 0 0
One HFRD 3 2.60 2 2 1 0
Two or more HFRD 5 4.04 4 3 3 0
Sedation procedure 7 4.04 4 3 3 0
Requires assistance 2 6.84 11 13 5 8
Unsteady gait 2 6.76 13 10 0 7
Visual/auditory impairment 2 6.92 6 2 9 11
Total Coefficient Sum 49 81.92 100 121 71 89
All optimized models have total coefficient sums greater than the current JHFRAT, and thus we compare feature contributions via their percent contributions to the total coefficient sum. A few trends persist across the model variations. Cognition risk factors feature far more prominently in the MIP models than in the current JHFRAT. Conversely, the history of falls risk factor only persists as non-zero in the asymmetric MIP model, albeit with less proportional importance (7.4%) compared to JHFRAT (10.2%).
4.2 Score Distributions and Categorization
Figure 2 reveals how different optimization objectives reshape the risk score distributions compared to the original JHFRAT. The Kolmogorov-Smirnov (KS) statistic quantifies the maximum vertical separation between the cumulative distribution functions of the Safe Low-Risk and Safe High-Risk cohorts, with larger values indicating better discriminative ability. The associated p-values test the null hypothesis that the two risk groups come from the same distribution—highly significant p-values (all p < 0.001) confirm that all models successfully separate the risk groups, though with varying degrees of separation.
Figure 2: Score distributions for all labeled data (test and train data together)
The distribution patterns directly reflect each model’s clinical optimization strategy:
Original JHFRAT–Conservative but Imprecise: The overlapping distributions (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage). This overlap drives both missed falls and unnecessary interventions. With 44.8% of eventual fall patients classified as low or moderate risk, critical prevention opportunities are missed.
Symmetric MIP–Maximum Separation for Clear Decision-Making: The stark bimodal distribution (KS = 0.699) eliminates the ambiguous middle ground, pushing patients toward clear low-risk or high-risk classifications. This suits healthcare systems that prefer definitive decisions: either minimal monitoring (score < 6) or full prevention protocols (score > 13). The trade-off is extreme scores (up to 60) that may seem implausible to clinicians, potentially undermining trust. Best for: facilities with binary intervention protocols where middle-ground treatments offer little value.
Asymmetric MIP–Aggressive Fall Prevention: The rightward shift of both distributions reflects the 3:1 penalty for under-triage, effectively lowering the bar for high-risk classification. This achieves 86% sensitivity for falls (vs. 45% in original JHFRAT) but classifies 17% of safe low-risk patients as high-risk. The extreme score range (up to 70) results from the optimizer pushing borderline cases strongly rightward to avoid missing falls. Best for: facilities where fall consequences are severe (e.g., surgical or elderly units) and resources allow broader intervention deployment.
FP-cap MIP–Targeted High-Risk Identification: By constraining false positives to current levels, this model (KS = 0.671) maintains clinician workload while better identifying true high-risk patients. The distribution shows selective rightward movement—pushing true fall patients across the high-risk threshold while keeping most low-risk patients in place. Scores remain clinically plausible (maximum 35). This achieves 45% fall sensitivity with 81% positive predictive value, meaning when it flags high-risk, it’s usually correct. Best for: resource-constrained settings where intervention capacity is fixed and false alarms cause alert fatigue.
FP&FN-cap MIP—Balanced Improvement Within Constraints: The bidirectional constraints produce the most JHFRAT-like distribution while still achieving better separation (KS = 0.683 vs. 0.620). Both cohorts shift toward the extremes but within bounds that maintain operational compatibility. This model essentially "tune” JHFRAT rather than replacing it, achieving 64% fall sensitivity while keeping false positive rates manageable. Best for: conservative healthcare systems requiring evidence that new models won’t disrupt operations before accepting larger changes.
CS–Smooth Risk Stratification: The continuous optimization produces graduated separation without the stark bimodality of MIP solutions. The smooth distribution (KS = 0.700) suggests CSO finds a natural risk continuum rather than forcing binary separation. This may better reflect clinical reality where risk truly is continuous, not discrete. With 66% fall sensitivity and moderate false positives, it balances competing objectives. Best for: initial implementations where extreme changes might face resistance, or as a warm-start for subsequent MIP refinement.
The key insight is that no single distribution is universally "best"–each reflects different clinical priorities and operational constraints. The original JHFRAT’s overlapping distributions explain its poor performance, while our optimized models offer a menu of alternatives: maximum separation (Symmetric MIP) for clear decisions, protective bias (Asymmetric MIP) for high-stakes environments, resource-aware precision (FP-cap) for efficiency, or incremental improvement (FP&FN-cap) for conservative adoption. The extreme scores in unconstrained models, while mathematically optimal, highlight the importance of including clinical face-validity constraints in future deployments.
We calculate the score differentials (MIP score - JHFRAT score) for all encounters in the dataset, as seen in Figure 3. All MIP models results in a positive average score differential, with a heavy upper tail. The asymmetric MIP without any performance constraints results in the most positive skew and the FP-cap MIP results in the least positive skew.
Figure 3: Score differentials (MIP score - JHFRAT score) for all patients in dataset
Skewness in the score differentials is also reflected in the categorization of encounters across the low, moderate and high-risk categories shown in Table 3. Without additional performance constraints, both the symmetric and asymmetric models significantly increase the number of encounters categorized as high-risk in the test data set. The symmetric model results in a 1100% increase in high-risk false positives and a 164% increase in true positives while the asymmetric model has a 1878% increase in high-risk false positives and 200% increase in true positives in the test set. The two models with false positive maximums still achieve notable gains in true positives: 56% and 124% increases from JHFRAT for FP-cap and FN&FP-cap, respectively. Both models also decrease the number of safe high-risk patients categorized as moderate-risk, but increase the number categorized as low-risk. All models except for the asymmetric MIP without performance constraints achieve a small increase in the number of true negative safe low-risk patients identified as low-risk.
Table 3: Test data encounter categorization across baseline and optimized models
Model Encounters Model Categorization
Low-Risk Moderate-Risk High-Risk
Original JHFRAT Safe Low-Risk 1,149 (72.6%) 420 (26.5%) 14 (0.9%)
Safe High-Risk 7 (8.1%) 55 (63.2%) 25 (28.7%)
All 17,007 (31.4%) 30,254 (55.8%) 6,946 (12.8%)
CSO Safe Low-Risk 1,156 (73.0%) 374 (20.5%) 53 (3.3%)
Safe High-Risk 8 (9.2%) 22 (25.3%) 57 (65.5%)
All 15,640 (28.9%) 18,124 (33.43%) 20,443 (37.7%)
Symmetric MIP Safe Low-Risk 1,180 (74.5%) 262 (16.6%) 168 (10.6%)
Safe High-Risk 7 (8.1%) 14 (16.1%) 66 (75.9%)
All 16,207 (29.9%) 10,581 (19.5%) 27,419 (50.58%)
Asymmetric MIP Safe Low-Risk 946 (59.8%) 360 (22.7%) 277 (17.5%)
Safe High-Risk 6 (6.9%) 6 (6.9%) 75 (86.2%)
All 11,278 (20.8%) 9,212 (17.0%) 33,717 (62.2%)
FP-cap MIP Safe Low-Risk 1,258 (79.5%) 316 (20.0%) 9 (0.6%)
Safe High-Risk 12 (13.8%) 36 (41.4%) 39 (44.8%)
All 18,142 (33.5%) 22,892 (42.2%) 13,172 (24.30%)
FP&FN-cap MIP Safe Low-Risk 1,237 (78.1%) 294 (18.6%) 52 (3.3%)
Safe High-Risk 11 (12.6%) 20 (23.0%) 56 (64.4%)
All 18,034 (33.3%) 15,562 (28.7%) 20,611 (38.0%)
4.3 Performance Metrics
Table 4 includes the performance metrics calculated for each model. In general, the optimized models achieve moderate gains in tight accuracy at the expense of loose accuracy. Only the FP-cap MIP improves on the tight accuracy without loss of loose accuracy.
Table 4: Performance metrics on test dataset
Model Accuracy AUC High-Risk Precision High-Risk Recall
Tight Loose AUROC AUPRC
Original JHFRAT 0.70 0.99 0.92 0.51 0.64 0.29
CSO 0.73 0.96 0.94 0.65 0.52 0.66
Symmetric MIP 0.75 0.90 0.94 0.65 0.28 0.76
Asymmetric MIP 0.61 0.83 0.92 0.57 0.21 0.86
FP-cap MIP 0.78 0.99 0.92 0.62 0.81 0.45
FN&FP-cap MIP 0.77 0.96 0.90 0.63 0.52 0.64
All optimization models outperform JHFRAT in area under the precision-recall curve for the testing and training datasets. All models perform similarly with high recall in the testing dataset, with the curves only notably separating at recall below 0.6. In contrast, the optimized models achieve visibly higher precision across all recall points for the training data. The small improvements in AUROC achieved by the models in the training set do not materialize in the testing set.
Figure 4: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves across JHFRAT and five optimization models on training dataset (above) and testing dataset (below).
5 Discussion
5.1 Key Findings
The proposed CSO and mixed-integer programming pipeline achives significant performance gains compared to the current JHFRAT model while offering priority-based tuning options in the hyperparameters. In general, the score distributions and sample categorizations respond as expected to the provided hyperparameters. Utilizing a symmetric distance function for mis-classification loss results in nearly symmetrical mis-classification rates for low-risk and high-risk labeled samples, while the use of an asymmetric 3-to-1 high-risk to low-risk distance function results in higher skew of both classes towards high-risk categorization. The notable exception to expectations is the increase in low-risk false negatives among the two performance-constrained models. The false-negative constraint for the FN&FP-cap MIP should mitigate an increase in such classifications, but this cannot be guaranteed in the testing setting.
While the CSO and MIP models themselves do not inherently discourage moderate-risk classification, when applied in a setting with no such moderate risk labels, the resulting score distributions disaggregate from the mean. This movement away from moderate risk classification can come at the expense of increasing further-class mis-classification, as is the case for the FP-cap and FN&FP-cap models. Nevertheless, these models remain promising options for improving true high-risk classification without drastic increases in the total number of patients classified as high-risk.
5.2 Limitations and Future Work
Several model extensions warrant investigation:
  • Non-linear scoring functions: Extending to tree-based or shallow neural architectures while maintaining interpretability
  • Time-varying risk: Incorporating temporal dynamics in sequential assessment settings
  • Multi-site learning: Federated optimization across institutions with heterogeneous populations
There are a few key limitations with the JHFRAT model experiments. Use of average JHFRAT scores over the encounter for model training results in some non-binary risk score values. These binary risk-score values represent averages over the encounter, but may contribute to extreme values in the optimized coefficients (i.e. coefficients larger than the highest class lower bound) in the absence of enforced upper bounds. The exclusion of all intervention cohort encounters also potentially limits the generalizability of the optimized models. Future work could explore including some patients from the intervention cohort as Safe Moderate-Risk, or apply multiple labels, depending on the number and type of fall prevention interventions applied.
5.3 Clinical Deployment Considerations
It is of particular clinical interest to improve identification of patients at high risk for falls while avoiding an overall increase in risk scores. From a resource perspective, it is impractical and wasteful to provide fall prevention interventions to patients who do not need them. Furthermore, from a clinical perspective, restricting mobility is an effective fall prevention method, but often leads to detrimental functional decline. Therefore, placing such restrictions on low-risk patients is not only unnecessary, but is actively harmful. These types of concerns over false positives extend beyond fall risk assessment. Alert fatigue is a persistent problem that is exacerbated by many false positives in assessment of risks of adverse events such as sepsis. The inclusion of performance constraints and the user-tunable distance function for mis-classification loss serve as tools to control the balance between under-classification and over-classification. Specific constraints on the number of false positives and/or false negatives for the most extreme classes offer assuredness.
In general, the specific features of the proposed MIP formulation that are applied to any given setting can be based on the context of the risk factors, adverse event in question, available data, and current clinical practice. For example, enforcing optimized coefficients to be integers may be unnecessary if assessments are completed automatically and clinicians only observe the final score. Allowing negativity in model coefficients allows for the inclusion of risk-preventative factors in the model.
6 Conclusion
In this work we develop a multi-class mixed integer programming ordinal classification model and a smooth, convex constrained score optimization. These models can apply in a variety of risk assessment contexts, and are particularly well suited for settings with uncertain ground truth or predictive performance requirements compared to an existing model. We utilize the proposed MIP and CSO formulations to conduct data-driven score optimization of the Johns Hopkins Fall Risk Assessment Tool. Across different model variations, we are able to increase identification of patients who fell as high-risk, and control for false positives via model parameters and constraints. Future work can explore refining or adapting the proposed models to other risk assessment settings.